AntonyTheGreatest:

what is the answer for polynomial 6x2 + x − 15 has a factor of 2x − 3. What is the other factor?

AntonyTheGreatest:

i am really confused

snowflake0531:

It wants you to find x where \(x(2x-3)=6x^2+x-15\) So if you know how to polynomial division or synthetic division, you could do \[\frac{6x^2+x-15}{2x-3}\]

AntonyTheGreatest:

these are the answer to choose from3x − 5 3x + 5 4x − 5 4x + 5

SmokeyBrown:

You're basically trying to find another expression that multiplies with (2x-3) in order to equal 6x^2 + x - 15 Since the first part of the polynomial is 6x^2, we know that the other factor must start with 3x, since 3x * 2x = 6x^2 What about the rest of the factor? What multiplies with -3 to get -15?

AntonyTheGreatest:

yea i have a lot of questions